# Euler-type approximation for the invariant measure: An abstract framework

### v1

@@ 1. Introduction
The problem of existence and uniqueness of the invariant measure for a Markov semigroup has a long history. We refer the reader to the influential paper by Meyn and Tweedie [MT93] and Hairer’s lecture notes [Hai21]. In direct connection to this question, the speed of convergence of the semigroup to the invariant measure has been also widely studied, see in particular Bakry et al. [BCG08], Cattiaux Guillin [CG09], Hairer [Hai21] and the references therein. Unless the semigroup can be simulated exactly, one typically uses approximation schemes of the semigroup to then approximate the invariant measure. Approximation schemes for the invariant measure associated to some stochastic process are thus built in two steps: a first step deals with the error between the invariant measure and the distribution of the semigroup for large times, and second step deals with the error between the semigroup and its approximation.
@@
In this sense, approximations based on Euler type schemes has also been an important research topic in the past years, particularly in the specific framework of the Langevin diffusion that is related to gradient descent algorithms that are widely used in Statistics and Machine Learning. We may distinguish two streams of research. The first one considers a constant time step for the Euler type approximation and then analyses the error between the invariant measures of the continuous and discretised semigroup, see in particular Talay [Tal02], Mattingly et al. [MSH02], Dalalyan and Tsybakov [DT12], Dalalyan [Dal17], Durmus and Moulines [DM17], Brosse et al. [BDMS19], Crisan et al. [CDO21] and Chen et al. [CDSX23]. The second stream of research, initiated by Lamberton and Pagès [LP02, LP03] and Lemaire [Lem05] considers instead an approximation scheme with a decreasing sequence of time steps, which allows to get directly the convergence to the invariant measure of the continuous semigroup. This point of view has been used in several recent works: see Pagès and Panloup [PP23], Bally and Qin [BQ24].
@@
Furthermore a number of stochastic equations in which the law of the solution is part of the equation, such as in the cases of McKean-Vlasov or Boltzmann type equations, do not seem to be within the framework of Markov semigroups. The literature about the approximation of the invariant measure for such systems is less developed but very active in the last years (see Schuh ans Souttar [SS24], Cormier [Cor24] and Du et. al. [DJL23] and the references therein). Having this motivation in mind, we focus on the approximations to the invariant measures for flows of transformations on the Wasserstein space, indexed on time. First, under appropriate hypotheses, we obtain exponential speed of convergence to the invariant measure as the time goes to infinity. Second, considering an approximation scheme with a decreasing step, we obtain in a general framework the rate of convergence in Wasserstein distance between the invariant measure and its approximation. Our work is related to the recent paper of Schuh and Souttar [SS24] who provide also a general framework to get uniform in time estimates between the continuous time process and its approximation, but for a constant time step approximation. As a consequence, there are significant differences between their framework and ours that we present in detail in Remark 2.13 below.
@@
with the supremum taken on all the partitions $\pi$ and all $n\in{\mathbb{N}}$. In many examples, this property is obtained with a Foster Lyapunov type condition (3.2). In some other examples, this property can be derived from existing results on the corresponding continuous time process.
@@
for 0<ts<h0<t-s<h. Lemma 2.6 proves then under this condition that
$W_{p}^{p}(\Theta_{t_{0},t_{n}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{n}}^{2,\pi}(\mu^{2}))\leq e^{-b_{*}(t_{n}-t_{0})}W_{p}^{p}(\mu^{1},\mu^{2})+C(1+M_{p}(\Theta^{1},\mu^{1},h)+\left\|\mu^{2}\right\|_{p}^{p})\sigma_{b_{*},\varepsilon}(n)$ (1.2)
with
$\sigma_{b_{*},\varepsilon}(n)=\sum_{k=1}^{n}e^{-b_{*}(t_{n}-t_{k})}\gamma_{k}^{1+\varepsilon}\quad\text{ and }\quad\ \gamma_{k}:=t_{k+1}-t_{k}.$
Following Lamberton and Pagès in [LP02] and [LP03], we get under suitable assumptions on the decreasing time step $\gamma_{n}$ that $\sigma_{b_{*},\varepsilon}(n)\to_{n\to 0}0$ with a precise rate of convergence. A typical example is to take $\gamma_{n}=\frac{1}{h^{-1}+n}$, and we have then $\sigma_{b_{*},\varepsilon}(n)\leq C/n^{\varepsilon}$.
We are now able to present our main abstract results [^1] For exact definitions in the results below, we refer the reader to the main text.. The first one concerns existence and uniqueness of the invariant measure and exponential estimates of the speed of convergence. Let $\theta_{s,t}:\mathcal{P}_{p}(B)\rightarrow\mathcal{P}_{p}(B),s\leq t$ be a flow ($\theta_{r,t}\circ\theta_{s,r}=\theta_{s,t}$ for $0\leq s\leq r\leq t$) that is time homogeneous (i.e. $\theta_{s,t}(\mu)=\theta_{0,t-s}(\mu)$) such that $\mu\rightarrow\theta_{s,t}(\mu)$ is continuous for the $W_{p}$ distance. We assume that $\theta$ has a uniform bound on the $p$ moments and moreover that $\theta$ is $(p,b_{*},\varepsilon)$-coupled with itself, meaning that (1.1) holds with $\Theta^{1}=\Theta^{2}=\theta$. Then, there exists a unique invariant probability measure $\nu$ (that is $\theta_{s,t}(\nu)=\nu$ for every $s\leq t$, and one has the following exponential speed of convergence
$\forall\mu\in\mathcal{P}_{p}(B),\quad W_{p}^{p}(\theta_{s,t}(\mu),\nu)\leq Ce^{-(b_{*}\wedge\varepsilon)(t-s)}$ (1.3)
when $b_{*}\not=\varepsilon$ (an extra multiplicative factor (1+(ts))(1+(t-s)) appears when $b_{*}=\varepsilon$). Our second result concerns the approximation of $\nu$ by means of an approximation scheme for the grid $\pi$ with $\gamma_{k}=\frac{1}{h^{-1}+k}$, $k\in\mathbb{N}$. If we suppose in addition to the previous hypotheses that $\Theta_{s,t}:\mathcal{P}_{p}(B)\rightarrow\mathcal{P}_{p}(B),s\leq t$ which is $(p,b_{*},\varepsilon)$-coupled with $\theta$ in the sense of (1.1), then for every $\mu\in\mathcal{P}_{p}(B)$
@@
In the second part of the paper, we present examples that fit our general framework. The key is to prove the $(p,b_{*},\varepsilon)$ - coupling property (1.2) between the approximation family $\Theta$ and the limit flow $\theta$. In most cases, this is done by using a global contraction property on the coefficients as in Section 5. However, in some specific examples related to the Langevin equation, Eberle [Ebe16] (see also [Sch24]) succeeds in proving the existence of the invariant measure and exponential convergence to equilibrium under a much weaker assumption where the contraction with respect to the Euclidean norm has to be in force outside a compact set only. The argument is based on the construction of a specific distance, which is equivalent to the Euclidean distance, under which strict contraction holds on the whole space. On [Ebe16, Example 1], we prove under the same conditions as the ones stated in [Ebe16] that the Euler scheme converges, and we give as well a rate of convergence. In fact, the formalism developed in our paper allows to use the estimates obtained in [Ebe16] for the flow in order to study the Euler approximation.
Now, we give the main structure of the article. The abstract framework is presented in Section 2 together with our main general results. Section 3 provides interesting tools to fit our general framework. Namely, we show how the sewing lemma can help to get the coupling property (1.2) for $W_{1}$ type distances, and give a Foster-Lyapunov type criterion to get uniform bounds on moments. Then, we give in Section 4 the results concerning the Langevin equation mentioned above. The following Section 5 is devoted to some more examples. In Subsection 5.1, we consider a stochastic equation with reflection driven by a Brownian motion with a $W_{2}$ distance. In Subsection 5.2, we deal with a recent neuronal model which has been introduced in [DMGLP15] and [FL16] and has recently been discussed in [CTV20]. This is a mean field type model driven by a Poisson Point process. We give two results, in terms of $W_{1}$ and $W_{2}$ using different global contraction assumptions in each case. In Subsection 5.3, we consider a Boltzmann type equation - that is an equation driven by a Poisson Point process with intensity depending on the law of the solution. Such equations appear in models of interacting particle systems. We consider two possible regimes: a finite variation regime with the distance $W_{1}$ and a martingale regime with the distance $W_{2}$. The further analysis of the particle systems or of the approximation based on the ergodic measure arising from these law dependent examples will be done in a further research. As many examples consist in doing verification of the required hypotheses of our general framework, they are often similar in nature, and we have put some of them in the additional material in Appendix B, keeping the most illustrative proofs in the main text.
@@ 2. Framework and main results
By the triangle inequality, we have $d(y,x_{0})^{p}\leq 2^{p-1}(d(x,x_{0})^{p}+d(x,y)^{p})$ and get for any $\rho\in\Pi(\mu,\nu)$
@@
Following the usual approach of decreasing steps for approximating the invariant measure we also assume that $(\gamma_{n})$ is a nonincreasing sequence for $n\in\mathbb{N}^{\ast}$. If $\Theta_{s,t}(\mu)$ represents the law of a one step Euler scheme with initial distribution $\mu$, then $\Theta_{t_{0},t_{i}}^{\pi}(\mu)$ is the distribution of the Euler scheme on the grid $\pi$ at the final time $t_{i}$ with initial law $\mu$.
@@
In the sequel we will use the notation, for $\delta>0$,
$M_{p}(\Theta,\mu,\delta)=\sup_{\pi:t_{i}-t_{i-1}\leq\delta}\sup_{n}\left\|\Theta_{t_{0},t_{n}}^{\pi}(\mu)\right\|_{p}^{p}$
with the supremum taken on all the partitions $\pi$ and all $n\in{\mathbb{N}}.$ The proof of the fact that $M_{p}(\Theta,\mu,\delta)$ is finite is not trivial. So, in Subsection 3.2 we give a criterion (of Foster Lyapunov type) which allows to prove it in several concrete examples. More precisely, if this criterion is verified, then (see (3.6))
$M_{p}(\Theta,\mu,\delta)\leq\left\|\mu\right\|_{p}^{p}+C.$ (2.4)
However there are other examples (see the section on the Langevin equation) where this result is already given and consequently we do not need the Foster Lyapunov criterion.
Definition 2.2.
2.1. Coupling
@@
The really important constants in the above definition are $b_{\ast}$ and $\varepsilon>0$ which appear in (2.5). They determine the speed of convergence in our approximation results, see (2.18) and (2.20). In contrast, the constant $C_{\ast}$ only influences the constants multiplying the speed on convergence in (2.18) and (2.20), and in this sense is less important.
In the sequel we will consider a decreasing sequence $\gamma_{n}\in(0,1),n\in{\mathbb{N}}$ such that $\sum_{n\geq 1}\gamma_{n}=\infty$ and we denote

### v2

# arXiv:2509.03971v2[math.PR] 02 Mar 2026

@@ 1. Introduction
The problem of existence and uniqueness of the invariant measure for a Markov semigroup has a long history. We refer the reader to the influential paper by Meyn and Tweedie [MT93] and Hairer’s lecture notes [Hai21]. In direct connection to this question, the speed of convergence of the semigroup to the invariant measure has been also widely studied, see in particular Bakry et al. [BCG08], Cattiaux and Guillin [CG09], Hairer [Hai21] and the references therein. Unless the semigroup can be simulated exactly, one typically uses approximation schemes of the semigroup to then approximate the invariant measure. Approximation schemes for the invariant measure associated to some stochastic process are thus built in two steps: a first step deals with the error between the invariant measure and the distribution of the semigroup for large times, and second step deals with the error between the semigroup and its approximation.
@@
In this sense, approximations based on Euler type schemes has also been an important research topic in the past years, particularly in the specific framework of the Langevin diffusion that is related to gradient descent algorithms that are widely used in Statistics and Machine Learning. We may distinguish two streams of research. The first one considers a constant time step for the Euler type approximation and then analyses the error between the invariant measures of the continuous and discretised semigroup, see in particular Talay [Tal02], Mattingly et al. [MSH02], Dalalyan and Tsybakov [DT12], Dalalyan [Dal17], Durmus and Moulines [DM17], Brosse et al. [BDMS19], Crisan et al. [CDO21] and Chen et al. [CDSX23]. The second stream of research, initiated by Lamberton and Pagès [LP02, LP03] and Lemaire [Lem05] considers instead an approximation scheme with a decreasing sequence of time steps, which allows to get directly the convergence to the invariant measure of the continuous semigroup. This point of view has been used in several recent works: see Pagès and Panloup [PP23], Bally and Qin [BQ24] and Chen et al.  [CJXX25].
@@
Furthermore a number of stochastic equations in which the law of the solution is part of the equation, such as in the cases of McKean-Vlasov or Boltzmann type equations, do not seem to be within the framework of Markov semigroups. The literature about the approximation of the invariant measure for such systems is less developed but very active in the last years (see Schuh and Souttar [SS24], Cormier [Cor25] and Du et. al. [DJL23] and the references therein). Having this motivation in mind, we focus on the approximations to the invariant measures for flows of transformations on the Wasserstein space, indexed on time. First, under appropriate hypotheses, we obtain exponential speed of convergence to the invariant measure as the time goes to infinity. Second, considering an approximation scheme with a decreasing step, we obtain in a general framework the rate of convergence in Wasserstein distance between the invariant measure and its approximation. Our work is related to the recent paper of Schuh and Souttar [SS24] who provide also a general framework to get uniform in time estimates between the continuous time process and its approximation, but for a constant time step approximation. As a consequence, there are significant differences between their framework and ours that we present in detail in Remark 2.17 below.
@@
with the supremum taken on all the partitions $\pi$ with decreasing time steps and all $n\in{\mathbb{N}}$. In many examples, this property is obtained with a Foster Lyapunov type condition (3.2). In some other examples, this property can be derived from existing results on the corresponding continuous time process.
@@
for 0<ts<h0<t-s<h. Let $t_{0}\geq 0$ and $t_{n}=t_{0}+\sum_{k=1}^{n}\frac{1}{1+k}$. Then, Corollary 2.9 proves under (1.1) that for any $\mu^{1},\mu^{2}\in\mathcal{P}_{p}(B)$, there exists $A\in{\mathbb{R}}_{+}$ such that
$W_{p}^{p}(\Theta_{t_{0},t_{n}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{n}}^{2,\pi}(\mu^{2}))\leq A\frac{1}{n^{b_{*}\wedge\varepsilon}},\ n\geq 0,$ (1.2)
up to a multiplicative logarithm term when $b_{*}=\varepsilon$.
Now we present the main consequences of the above general abstract result that will be proven throughout the article [^1] For exact definitions and conditions, we refer the reader to the main text.. The first one concerns existence and uniqueness of the invariant measure and exponential estimates of the speed of convergence. Let $\theta_{s,t}:\mathcal{P}_{p}(B)\rightarrow\mathcal{P}_{p}(B),s\leq t$ be a flow ($\theta_{r,t}\circ\theta_{s,r}=\theta_{s,t}$ for $0\leq s\leq r\leq t$) that is time homogeneous (i.e. $\theta_{s,t}(\mu)=\theta_{0,t-s}(\mu)$) such that $\mu\rightarrow\theta_{s,t}(\mu)$ is continuous for the $W_{p}$ distance. We assume that $\theta$ has a uniform bound on the $p$ moments and moreover that $\theta$ is $(p,b_{*},\varepsilon)$-coupled with itself, meaning that (1.1) holds with $\Theta^{1}=\Theta^{2}=\theta$. Then, there exists a unique invariant probability measure $\nu$ (that is $\theta_{s,t}(\nu)=\nu$ for every $s\leq t$, and one has the following exponential speed of convergence
$\forall\mu\in\mathcal{P}_{p}(B),\exists C\in\mathbb{R}_{+},\forall t\geq s,\ \quad W_{p}^{p}(\theta_{s,t}(\mu),\nu)\leq Ce^{-(b_{*}\wedge\varepsilon)(t-s)}$ (1.3)
when $b_{*}\not=\varepsilon$ (an extra multiplicative factor (1+(ts))(1+(t-s)) appears when $b_{*}=\varepsilon$). Our second result concerns the approximation of $\nu$ by means of an approximation scheme for the grid $\pi$ with $\gamma_{k}=\frac{1}{1+k}$, $k\in\mathbb{N}$. If we suppose in addition to the previous hypotheses that $\Theta_{s,t}:\mathcal{P}_{p}(B)\rightarrow\mathcal{P}_{p}(B),s\leq t$ which is $(p,b_{*},\varepsilon)$-coupled with $\theta$ in the sense of (1.1), then for every $\mu\in\mathcal{P}_{p}(B)$, there exists $C\in\mathbb{R}_{+}$ such that
@@
In the second part of the paper, we present examples that fit our general framework. The key is to prove the $(p,b_{*},\varepsilon)$ - coupling property (1.2) between the approximation family $\Theta$ and the limit flow $\theta$. In most cases, this is done by using a global contraction property on the coefficients and a short time analysis as in Section 5. However, in some specific examples related to the Langevin equation, Eberle [Ebe16] (see also [Sch24]) succeeds in proving the existence of the invariant measure and exponential convergence to equilibrium under a much weaker assumption where the contraction with respect to the Euclidean norm has to be in force outside a compact set only. The argument is based on the construction of a specific distance, which is equivalent to the Euclidean distance, under which strict contraction holds on the whole space. On [Ebe16, Example 1], we prove under the same conditions as the ones stated in [Ebe16] that the Euler scheme converges, and we give as well a rate of convergence. In fact, the formalism developed in our paper allows to use the estimates obtained in [Ebe16] for the flow in order to study the Euler approximation.
Now, we give the main structure of the article. The abstract framework is presented in Section 2 together with our main general results. Section 3 provides interesting tools to fit our general framework. Namely, we show how the sewing lemma can help to get the coupling property (1.2) for $W_{1}$ type distances, and give a Foster-Lyapunov type criterion to get uniform bounds on moments. Then, we give in Section 4 the results concerning the Langevin equation mentioned above. The following Section 5 is devoted to some more examples. In Subsection 5.1, we consider a stochastic equation with reflection driven by a Brownian motion with a $W_{2}$ distance. In Subsection 5.2, we deal with a recent neuronal model which has been introduced in [DMGLP15] and [FL16] and has recently been discussed in [CTV20]. This is a mean field type model driven by a Poisson Point process. We give two results, in terms of $W_{1}$ and $W_{2}$ using different global contraction assumptions in each case. In Subsection 5.3, we consider a Boltzmann type equation - that is an equation driven by a Poisson Point process with intensity depending on the law of the solution. Such equations appear in models of interacting particle systems. We consider two possible regimes: a finite variation regime with the distance $W_{1}$ and a martingale regime with the distance $W_{2}$. The further analysis of the particle systems or of the approximation based on the ergodic measure arising from these law dependent examples will be done in a further research. As many examples consist in doing verification of the required hypotheses of our general framework, they are often similar in nature, and we have put some of them in the additional material in Appendix C, keeping the most illustrative proofs in the main text.
@@ 2. Framework and main results
It will be used as a shorthand notation in different upper bounds. By the triangle inequality, we have $d(y,x_{0})^{p}\leq 2^{p-1}(d(x,x_{0})^{p}+d(x,y)^{p})$ and get for any $\rho\in\Pi(\mu,\nu)$
@@
Following the usual approach of decreasing steps for approximating the invariant measure we also assume that $(\gamma_{n})_{n\in\mathbb{N}^{\ast}}$ is a nonincreasing sequence. If $\Theta_{s,t}(\mu)$ represents the law of a one step Euler scheme with initial distribution $\mu$, then $\Theta_{t_{0},t_{i}}^{\pi}(\mu)$ is the distribution of the Euler scheme on the grid $\pi$ at the final time $t_{i}$ with initial law $\mu$. If $\Theta_{s,t}$ enjoys the flow property $\Theta_{s,t}(\Theta_{r,s}(\mu))=\Theta_{r,t}(\mu)$ for $r\leq s\leq t$, then $\Theta_{t_{0},t_{i}}^{\pi}(\mu)=\Theta_{t_{0},t_{i}}(\mu)$.
@@
In the case where $\Theta^{1}$ is the flow associated to a continuous time process and $\Theta^{2}$ is an approximation of $\Theta^{1}$, the constant $b_{*}$ can be thought as the contraction behaviour of the process, while $\varepsilon$ is related to the accuracy of the approximation in short time.
For technical reasons, we will need upper bounds on $p$-moments, and we introduce the notation, for $\delta>0$,
$M_{p}(\Theta,\mu,\delta)=\sup_{\|\pi\|<\delta}\sup_{n}\left\|\Theta_{t_{0},t_{n}}^{\pi}(\mu)\right\|_{p}^{p},$
with the supremum taken on all the partitions $\pi$ such that $(t_{i+1}-t_{i})_{i\geq 0}$ is decreasing and all $n\in{\mathbb{N}}$.
Definition 2.3.
The fact that $M_{p}(\Theta,\mu,\delta)$ is finite is not trivial. So, in Subsection 3.2 we give a Foster Lyapunov type criterion which allows to prove it in several concrete examples. However, there are other examples (see Remark 4.2) where this result is already given and consequently we do not need the Foster Lyapunov criterion.
We are now in position to state the main result of our paper.
Theorem 2.4.
Let $(\theta_{s,t})_{0\leq s\leq t}$ be a family of operators on $\mathcal{P}_{p}(B)$ with the flow property ($\theta_{s,t}\circ\theta_{r,s}=\theta_{r,t}$ for $r\leq s\leq t$). We assume that $\theta$ is time homogeneous (i.e. $\theta_{s,t}=\theta_{0,t-s}$) and such that for all $s\leq t$, $\mu\mapsto\theta_{s,t}(\mu)$ is continuous with respect to $W_{p}$. We assume besides that $\theta$ is $p$-bounded and $(p,b_{*},\varepsilon)$ self coupled. Then, there exists a unique invariant measure $\nu\in\mathcal{P}_{p}(B)$ associated to $\theta$, and for every $\mu\in\mathcal{P}_{p}(B)$, there exists $A\in\mathbb{R}_{+}$ such that
$W_{p}^{p}(\theta_{s,t}(\mu),\nu)\leq Ae^{-\varepsilon\wedge b_{\ast}(t-s)}(1+(t-s)^{\mathbf{1}_{b_{\ast}=\varepsilon}}).$
Let $(\Theta_{s,t})_{0\leq s\leq t}$ be a family of operators on $\mathcal{P}_{p}(B)$ such that $\theta$ and $\Theta$ are $(p,b_{*},\varepsilon)$ coupled. Let $a\geq 0$, $\gamma_{n}=\frac{1}{a+n}$ and $t_{n}=s+\gamma_{1}+...+\gamma_{n}$, $n\in{\mathbb{N}}$. Then, for every $\mu\in\mathcal{P}_{p}(B)$, there exists $A\in{\mathbb{R}}_{+}$ such that
$W_{p}^{p}(\Theta_{s,t_{n}}^{\pi}(\mu),\nu)\leq An^{-b_{\ast}\wedge\varepsilon}(1+\ln(n+1)^{\mathbf{1}_{b_{\ast}=\varepsilon}}).$
Theorem 2.4 is a consequence of Proposition 2.14 and Theorem 2.15. Note that here, the constant $A$ may depend on $\mu$. Under further assumptions, we will be able to get $A\leq C(1+\|\mu\|_{p}^{p})$, with $C\in{\mathbb{R}}_{+}$ that does not depend on $\mu$.
Many examples that illustrate this theorem are given in Sections 4 and 5. Here, we presents quickly two illustrative examples. The first one, perhaps the simplest one, is the Ornstein-Uhlenbeck process
$X^{1}_{s,t}(X^{1})=X^{1}-\int_{s}^{t}kX^{1}_{s,u}(X^{1})du+\sigma(W_{t}-W_{s}),\ s\leq t,$ (2.5)
with $k,\sigma>0$ and $X^{1}$ being a $p$-integrable real random variable independent of $W$. For $\mu^{1}\in\mathcal{P}_{p}(\mathbb{R})$, we define $\Theta^{1}_{s,t}(\mu^{1})$ as the law of $X^{1}_{s,t}(X^{1})$ when $X^{1}\sim\mu^{1}$. For $\mu^{2}\in\mathcal{P}_{p}(\mathbb{R})$ and $X^{2}\sim\mu^{2}$ independent of $W$, we define $\Theta^{2}_{s,t}(\mu^{2})$ as the law of the Euler-Maruyama scheme $X^{2}_{s,t}(X^{2})$, with
$X^{2}_{s,t}(X^{2})=(1-k(t-s)X^{2})+\sigma(W_{t}-W_{s}),\ s\leq t.$ (2.6)
In Appendix B, we prove that $\Theta^{1}$ is $p$-bounded and $(p,b_{*}p,\varepsilon)$ self-coupled for any $\varepsilon>0$ and $b_{*}\in(0,k)$. Besides, $\Theta^{1}$ and $\Theta^{2}$ are $(p,b_{*}p,\varepsilon)$ coupled for any $\varepsilon\in(0,p)$ and $b_{*}\in(0,k)$. Therefore, Theorem 2.4 gives, for any $\mu\in\mathcal{P}_{p}(\mathbb{R})$ and $\zeta\in(0,1)$, the existence of $A\geq 0$ such that
$W_{p}^{p}(\theta_{s,t}(\mu),\nu)\leq Ae^{-pb_{*}(t-s)},\ W_{p}^{p}(\Theta^{\pi}_{s,t_{n}}(\mu),\nu)\leq An^{-\zeta p}.$
By explicit calculations, we show on the other hand that $W_{p}^{p}(\theta_{s,t}(\mu),\nu)=_{t\to\infty}O(e^{-k(t-s)})$ and $W_{p}^{p}(\Theta^{\pi}_{s,t_{n}}(\mu),\nu)=_{n\to\infty}O(n^{-{k\wedge 1}})$. Therefore, the rates of convergence given by Theorem 2.4 are almost optimal on this example.
The second example is the McKean Vlasov equation
$X^{1}_{s,t}=X^{1}+\int_{s}^{t}\beta(X^{1}_{s,u},\mathcal{L}(X^{1}_{s,u}))du+\int_{s}^{t}\sigma(X^{1}_{s,u},\mathcal{L}(X^{1}_{s,u}))dW_{u},\ s\leq t,$
with $\beta:\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}^{d}$, $\sigma:\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}^{d\times d}$ and $W$ is a $d$-dimensional Brownian motion. To keep simple notation, we assume $d=1$ even if what follows works for any $d\geq 1$. We make the following assumption: there exists $C,k,L\geq 0$ such that for all $x,y\in{\mathbb{R}}$ and $\mu^{1},\mu^{2}\in\mathcal{P}_{2}({\mathbb{R}})$,
  1. (1)
    $|\beta(x,\mu^{1})-\beta(y,\mu^{2})|^{2}+|\sigma(x,\mu^{1})-\sigma(y,\mu^{2})|^{2}\leq C(|x-y|^{2}+W_{2}^{2}(\mu^{1},\mu^{2}))$,
  2. (2)
    $2(\beta(x,\mu^{1})-\beta(y,\mu^{2}))(x-y)+|\sigma(x,\mu^{1})-\sigma(y,\mu^{2})|^{2}\leq LW_{2}^{2}(\mu^{1},\mu^{2})-k|x-y|^{2}$.
Then, [Wan18, Theorem 2.1] gives the strong existence and uniqueness, and we note $\Theta^{1}_{s,t}(\mu^{1})$ the law of $X^{1}_{s,t}$ when $X^{1}\sim\mu^{1}$. If $k>L$, [Wan18, Theorem 3.1] gives $W_{2}^{2}(\Theta^{1}_{s,t}(\mu^{1}),\Theta^{1}_{s,t}(\mu^{2}))\leq W_{2}^{2}(\mu^{1},\mu^{2})e^{-(k-L)(t-s)}$, and therefore $\Theta^{1}$ is $(2,b_{*},\varepsilon)$ self coupled for any $b_{*}\in(0,k-L)$ and $\varepsilon>0$.
Let $\Theta^{2}_{s,t}(\mu^{2})$ the law of
$X^{2}_{s,t}=X^{2}+\beta(X^{2},\mu^{2})(t-s)+\sigma(X^{2},\mu^{2})(W_{t}-W_{s})$
when $X^{2}\sim\mu^{2}$. By Itô formula and using (2), we get
$\mathbb{E}[(X^{1}_{s,t}-X^{2}_{s,t})^{2}]\leq$ $\mathbb{E}[(X^{1}-X^{2})^{2}]$
$+\int_{s}^{t}LW_{2}^{2}(\mathcal{L}(X^{1}_{s,u}),\mathcal{L}(X^{2}_{s,u}))-k\mathbb{E}[(X^{1}_{s,u}-X^{2}_{s,u})^{2}]du$
$+\int_{s}^{t}2\mathbb{E}[(\beta(X^{2}_{s,u},\mathcal{L}(X^{2}_{s,u}))-\beta(X^{2},\mu^{2}))(X^{1}_{s,u}-X^{2}_{s,u})]du$
$\hskip-85.35826pt+\int_{s}^{t}\mathbb{E}[(\sigma(X^{1}_{s,u},\mathcal{L}(X^{1}_{s,u}))-\sigma(X^{2},\mu^{2}))^{2}-(\sigma(X^{1}_{s,u},\mathcal{L}(X^{1}_{s,u}))-\sigma(X^{2}_{s,u},\mathcal{L}(X^{2}_{s,u})))^{2}]du$
We now use $W_{2}^{2}(\mathcal{L}(X^{1}_{s,u}),\mathcal{L}(X^{2}_{s,u}))\leq\mathbb{E}[(X^{1}_{s,u}-X^{2}_{s,u})^{2}]$ for the first term of the right hand side. For $\delta>0$ arbitrarily small, we use $2ab\leq\delta a^{2}+\delta^{-1}b^{2}$ and $(a+b)^{2}\leq a^{2}(1+\delta)+b^{2}(1+\delta^{-1})$ respectively for the second and third terms, together with the Lipschitz property of $b$ and $\sigma$, and we get
$\mathbb{E}[(X^{1}_{s,t}-X^{2}_{s,t})^{2}]\leq$ $\mathbb{E}[(X^{1}-X^{2})^{2}]+\int_{s}^{t}(L+\delta-k)\mathbb{E}[(X^{1}_{s,u}-X^{2}_{s,u})^{2}]du$
$+C_{\delta}\int_{s}^{t}\mathbb{E}[(X^{2}_{s,u}-X^{2})^{2}]du.$
By standard estimates, we have $\mathbb{E}[(X^{2}_{s,u}-X^{2})^{2}]\leq\Gamma_{2}(\mu^{2},\mu^{2})(u-s)^{1/2}$. By taking $X^{1}$ and $X^{2}$ optimally coupled for $W_{2}$ and using Gronwall lemma, we deduce that $\Theta^{1}$ and $\Theta^{2}$ are $(2,b_{*},1/2)$ coupled for any $b_{*}\in(0,k-L)$. We can thus apply Theorem 2.4 to $\theta=\Theta^{1}$ and $\Theta=\Theta^{2}$ and get
$W_{2}^{2}(\Theta_{s,t_{n}}^{\pi}(\mu),\nu)\leq An^{-b_{*}\wedge\frac{1}{2}},$ (2.7)
for any $b_{*}\in(0,1/2]\cap(0,k-L)$, where $t_{n}=s+\sum_{i=1}^{n}\frac{1}{1+i}$.
@@
As the above example shows, the most important constants that appear in Definition 2.2, Eq. (2.4), are $b_{\ast}$ and $\varepsilon>0$. In practical examples, $b_{*}$ corresponds to the mean-reverting behaviour of the process while $\varepsilon$ is mostly related to the short time approximation error of the chosen scheme (see also examples in Section 5). They determine the speed of convergence in our approximation results of the stationary distribution in Theorem 2.4. This will appear clearly in the proof of Lemma 2.6. In contrast, the constant $C_{\ast}$ only influences the constants multiplying the speed on convergence, and is in this sense is less important.
2.1. Coupling on a time grid
In the sequel we will consider a decreasing sequence $\gamma_{n}\in(0,1),n\in{\mathbb{N}}$ such that $\lim_{n\to\infty}\gamma_{n}=0$, $\sum_{n\geq 1}\gamma_{n}=\infty$ and we denote